A physical fitness association is including the mile run in its secondary-school fitness test. The
time for this event for boys in secondary school is known to possess a normal distribution with a
mean of 450 seconds and a standard deviation of 50 seconds. Find the probability that a randomly
selected boy in secondary school can run the mile in less than 335 seconds.
Answer #1
The mean is 450 seconds and the standard deviation is 50 seconds.
Number of standard deviations below mean
=(450-335)/50 = 2.3
Since it is assumed to be normally distributed, look up the normal distribution table to get the probability of 2.3 standard deviations lower than the mean. (approx. 1%)
answered by: deava
Answer #2

z= (335-450)/50

z= -115/50

z= -2.3 (0.0107)


1.07%

source: Math
answered by: DeMarcus